OpenStudy (anonymous):

Multiply 6 3/8 by 4/17.

OpenStudy (oaktree):

51/8 x 4/17 = 51/2 x 1/17 = 3/2

OpenStudy (anonymous):

Soo..

OpenStudy (anonymous):

start ny changing the first fraction to a top heavy by 6*8+3 =51 51/8 * 4/17 = 204/136 then simplify

OpenStudy (anonymous):

204/136 =1.5 or as OKTREE has it 3/2

OpenStudy (anonymous):

The answer is 1/1/2

OpenStudy (anonymous):